# CodeAssistBench (CAB): Dataset & Benchmarking for Multi-turn Chat-Based Code Assistance

### v4

@@ B.2 Inter-Rater Reliability
The two expert annotators demonstrated substantial overall agreement at 78.28% (728/930 judgments across all dimensions; Cohen’s $\kappa=0.68$), establishing a robust baseline for evaluating LLM judge performance. According to established guidelines landis1977measurement, $\kappa$ values between 0.61–0.80 indicate substantial agreement, confirming that our annotation protocol achieves a high level of consistency. Agreement varied across evaluation dimensions, as shown in Table 4.
@@
This variation reveals that while response style (verbosity) is relatively objective and unambiguous ($\kappa=0.77$, substantial agreement), assessing technical accuracy requires nuanced domain expertise and can yield different but equally defensible judgments ($\kappa=0.62\text{--}0.63$, moderate to substantial agreement). According to established guidelines landis1977measurement, $\kappa$ values between 0.61–0.80 indicate substantial agreement, and our results fall within this range across all dimensions. This level of agreement demonstrates that even experienced practitioners may weigh correctness, completeness, and stylistic factors differently when evaluating subjective technical tasks.
@@ B.3 LLM-Expert Agreement
B.4 Verified Dataset Release
Based on this validation, we release a verified subset of 149 issues (84.7% of the validation set) where at least one model response achieved full annotator concordance across all three evaluation dimensions. This verified subset provides high-quality ground truth for code assistance evaluation while maintaining broad coverage across programming languages and issue types: JavaScript (37), Python (37), TypeScript (23), C# (19), C++ (14), C (13), and Java (6). The strong verification rate across all languages validates the quality and clarity of our benchmark, providing researchers with reliable ground truth for benchmarking code assistance systems across diverse programming contexts. We plan to extend this verified subset in future work through additional expert validation.
@@ Appendix C Extended Discussion of Limitations
Finally, we do not report error bars or statistical significance tests. While aggregate model trends are clear, fine-grained comparisons—especially between close-performing models—should be interpreted cautiously.
@@ Appendix D Broader Impacts
CodeAssistBench (CAB) has the potential to advance the development of safer and more capable AI programming assistants, particularly in realistic, multi-turn developer scenarios. By surfacing failure modes and limitations in current models, CAB encourages the creation of tools that better align with developer needs, which may enhance productivity and learning. However, there are also potential negative impacts. For instance, if developers rely too heavily on inaccurate AI suggestions in real-world settings, this could introduce subtle bugs or security issues. Additionally, the use of scraped GitHub data—even when filtered—may raise concerns around privacy or attribution if not carefully managed. We encourage responsible use of CAB, including proper consent for dataset extension and monitoring downstream usage of models trained or benchmarked using this framework.
@@ Appendix I Human Annotation Instructions
We randomly selected 50 problems from the recent repository subset and generated fully synthetic reproductions using a maintainer agent backed by Claude Sonnet 4.5. For each issue, the agent was run iteratively until it confirmed the issue was fully reproducible within the generated codebase. Each synthetic repository maintained similar complexity (20-28 files, 1,500+ lines of code) while preserving the identical bugs and satisfaction conditions from the original issues. The synthetic repositories were designed to use programming language versions and library versions that fall within the model’s training knowledge cutoff.
@@
Table 8: Performance comparison across repository types
@@
    Training Data Recency Matters: The 57.51 percentage point difference between recent repositories (16.49%) and synthetic repositories (74.0%) primarily reflects knowledge cutoff limitations rather than fundamental reasoning deficits.
@@
To understand model strengths and weaknesses across different problem types, we performed automated error categorization on the 310 AI responses used in our expert validation study (Section B). This analysis clusters issues into coarse bug types, revealing targeted patterns in model capabilities.
@@
We used Claude Sonnet 4.5 to categorize each GitHub issue into one of seven error types based on the issue title, body, and conversation context. The categories are:
@@
Table 9: Distribution of error categories across 310 analyzed issues
@@
While CodeAssistBench focuses on conversational programming assistance, complementary work in automated API testing demonstrates the value of specification-grounded evaluation that shares conceptual similarities with our satisfaction condition approach. Recent research has explored automated testing for REST APIs using OpenAPI specifications as structured artifacts kim2022automated, kim2023enhancing, leveraging NLP techniques to extract semantic information including parameter constraints and endpoint dependencies. Subsequent work has demonstrated how large language models can improve REST API testing by generating semantically meaningful test inputs kim2024leveraging, with multi-agent systems further enhancing test effectiveness by modeling complex API interactions kim2024multi. More recent work explores whether small language models can effectively handle domain-specific testing tasks when provided with clear specifications kim2025llamaresttest, demonstrating that specification quality significantly impacts model effectiveness—a finding that aligns with our observation that documentation quality affects resolution success in CAB. While API testing benchmarks target functional correctness through executable specifications (OpenAPI) and CAB evaluates conversational assistance through extracted satisfaction conditions, both approaches share a fundamental principle: specification-grounded evaluation that moves beyond subjective assessment toward principled evaluation based on explicitly stated requirements, whether derived from formal specifications or real developer conversations.

### v5

@@ B.2 Inter-Rater Reliability
The two expert annotators demonstrated substantial overall agreement at 78.28% (728/930 judgments across all dimensions; Cohen’s $\kappa=0.68$), establishing a robust baseline for evaluating LLM judge performance. According to established guidelines Landis and Koch [1977], $\kappa$ values between 0.61–0.80 indicate substantial agreement, confirming that our annotation protocol achieves a high level of consistency. Agreement varied across evaluation dimensions, as shown in Table 4.
@@
This variation reveals that while response style (verbosity) is relatively objective and unambiguous ($\kappa=0.77$, substantial agreement), assessing technical accuracy requires nuanced domain expertise and can yield different but equally defensible judgments ($\kappa=0.62\text{--}0.63$, moderate to substantial agreement). According to established guidelines Landis and Koch [1977], $\kappa$ values between 0.61–0.80 indicate substantial agreement, and our results fall within this range across all dimensions. This level of agreement demonstrates that even experienced practitioners may weigh correctness, completeness, and stylistic factors differently when evaluating subjective technical tasks.
@@ Appendix C Extended Discussion of Limitations
Seventh, we do not report error bars or statistical significance tests. While aggregate model trends are clear, fine-grained comparisons—especially between close-performing models—should be interpreted cautiously.
Lastly, the 12.36 percentage point gap between our LLM judge’s agreement with human annotators (65.92%) and the inter-human baseline (78.28%) represents a fundamental limitation of automated multi-turn evaluation. This gap is particularly pronounced for subjective dimensions like technical correctness (54–65% LLM-human agreement) compared to objective dimensions like verbosity (74–79%).
We hypothesize that multi-turn conversations introduce several sources of variance that make alignment harder than single-turn evaluation:
  1. 1.
    Accumulated context interpretation: As conversations progress, the accumulated context can be interpreted differently by human and LLM judges. A response that seems adequate given early turns may be judged differently when later clarifications reveal the user’s true intent.
  2. 2.
    Credit attribution ambiguity: In multi-turn dialogues, it becomes unclear which turn’s contribution should be credited for ultimate success or failure. Human annotators may attribute success to the final clarifying response, while LLM judges may weight earlier turns more heavily.
  3. 3.
    Trajectory-dependent evaluation: The combinatorial space of possible conversation trajectories makes it difficult to establish consistent evaluation criteria. Two conversations reaching the same endpoint via different paths may receive different judgments depending on the intermediate steps.
  4. 4.
    Error propagation: Early misunderstandings or incomplete responses can cascade through subsequent turns, making it challenging to isolate which specific failure led to an incorrect final assessment.
This finding has important implications for the broader research community working on LLM-as-judge methods and multi-turn dialogue evaluation. We suggest several directions for future work to address this limitation:
  • Turn-level annotations: Rather than evaluating entire conversations holistically, decomposing evaluation into per-turn assessments may reduce variance and improve alignment.
  • Trajectory-based metrics: Developing metrics that explicitly account for conversation dynamics and progression, rather than treating dialogues as atomic units.
  • Consensus-based judging: Using multiple LLM judges with ensemble methods to reduce individual model biases, similar to how we used multiple human annotators.
  • Calibrated confidence scores: Training judges to output calibrated confidence scores that reflect uncertainty in multi-turn contexts, enabling downstream systems to weight judgments appropriately.
We leave the systematic investigation of these approaches to future work, but note that improving multi-turn evaluation reliability remains critical for advancing conversational code assistance research.
@@ Appendix D Broader Impacts
CodeAssistBench (CAB) has the potential to advance the development of safer and more capable AI programming assistants, particularly in realistic, multi-turn developer scenarios. By surfacing failure modes and limitations in current models, CAB encourages the creation of tools that better align with developer needs, which may enhance productivity and learning. However, there are also potential negative impacts. For instance, if developers rely too heavily on inaccurate agent suggestions in real-world settings, this could introduce subtle bugs or security issues. Additionally, the use of scraped GitHub data—even when filtered—may raise concerns around privacy or attribution if not carefully managed. We encourage responsible use of CAB, including proper consent for dataset extension and monitoring downstream usage of models trained or benchmarked using this framework.
@@ Appendix I Human Annotation Instructions
We randomly selected 50 problems from the recent repository subset and generated fully synthetic reproductions using a maintainer agent backed by Claude Sonnet 3.7. For each issue, the agent was run iteratively until it confirmed the issue was fully reproducible within the generated codebase. Each synthetic repository maintained similar complexity (20-28 files, 1,500+ lines of code) while preserving the identical bugs and satisfaction conditions from the original issues. The synthetic repositories were designed to use programming language versions and library versions that fall within the model’s training knowledge cutoff.
@@
Table 8: Performance comparison across repository types (Sonnet 3.7)
@@
    Training Data Recency Matters: The 62.66 percentage point difference between recent repositories (11.34%) and synthetic repositories (74.0%) primarily reflects knowledge cutoff limitations rather than fundamental reasoning deficits.
@@
To understand model strengths and weaknesses across different problem types, we performed automated error categorization on the 620 human-annotation records (310 agent responses $\times$ 2 annotators) from our expert validation study (Section B). This analysis clusters issues into coarse bug types, revealing targeted patterns in model capabilities.
@@
We used Claude Sonnet 3.7 to categorize each GitHub issue into one of seven error types based on the issue title, body, and conversation context. The categories are:
@@
Table 9: Distribution of error categories across 620 annotation records
@@
While CodeAssistBench focuses on conversational programming assistance, complementary work in automated API testing demonstrates the value of specification-grounded evaluation that shares conceptual similarities with our satisfaction condition approach. Recent research has explored automated testing for REST APIs using OpenAPI specifications as structured artifacts Kim et al. [2022, 2023], leveraging NLP techniques to extract semantic information including parameter constraints and endpoint dependencies. Subsequent work has demonstrated how large language models can improve REST API testing by generating semantically meaningful test inputs Kim et al. [2024], with multi-agent systems further enhancing test effectiveness by modeling complex API interactions Kim et al. [2025b]. More recent work explores whether small language models can effectively handle domain-specific testing tasks when provided with clear specifications Kim et al. [2025a], demonstrating that specification quality significantly impacts model effectiveness—a finding that aligns with our observation that documentation quality affects resolution success in CAB. While API testing benchmarks target functional correctness through executable specifications (OpenAPI) and CAB evaluates conversational assistance through extracted satisfaction conditions, both approaches share a fundamental principle: specification-grounded evaluation that moves beyond subjective assessment toward principled evaluation based on explicitly stated requirements, whether derived from formal specifications or real developer conversations.